OpenStudy (anonymous):

Someone please help D: I really suck at math can anyone please help. I'd appreciate it so much! Find the focus and directrix of the graph of the equation. Equation: y=5x^2 Thanks!

OpenStudy (jack1):

Equation: y=5x^2 Equation for focus: x^2 = 4ay so rearrange your equation ( y=5x^2) into this form (x^2 = 4ay), and the point (0,a) will be your focus

OpenStudy (jack1):

so if u start with your equation: y=5x^2 y/5=x^2 - divide both sides by 5 now u have x^2 = y/5 and we know x^2 = 4ay so 4a = 1/5 so a = 1/20 = 0.05

OpenStudy (jack1):

so your focus in this case will be at (0, a) which in this case is (0, 0.05)

OpenStudy (jack1):

directrix... never understood it, im sorry u could try http://www.purplemath.com/modules/parabola.htm or we could see if @ganeshie8 has some free time?

ganeshie8 (ganeshie8):